### v1

# Mini-InternVL: A Flexible-Transfer Pocket Multimodal Model with 5% Parameters and 90% Performance

Abstract
Multimodal large language models (MLLMs) have demonstrated impressive performance in vision-language tasks across a broad spectrum of domains. However, the large model scale and associated high computational costs pose significant challenges for training and deploying MLLMs on consumer-grade GPUs or edge devices, thereby hindering their widespread application. In this work, we introduce Mini-InternVL, a series of MLLMs with parameters ranging from 1B to 4B, which achieves 90% of the performance with only 5% of the parameters. This significant improvement in efficiency and effectiveness makes our models more accessible and applicable in various real-world scenarios. To further promote the adoption of our models, we develop a unified adaptation framework for Mini-InternVL, which enables our models to transfer and outperform specialized models in downstream tasks, including autonomous driving, medical images, and remote sensing. We believe that our study can provide valuable insights and resources to advance the development of efficient and effective MLLMs.
1 Introduction
In recent years, there have been significant advancements in multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6], which leverages the powerful capabilities of pre-trained large language models (LLMs) [7, 8, 9, 10, 11, 12] alongside vision foundation models (VFMs) [1, 13, 14]. These models undergo multi-stage training on extensive image-text data, which effectively aligns visual representations from VFMs with the latent space of LLMs, leading to promising performance in general vision-language understanding, reasoning, and interaction tasks. However, the large computational burden and the poor performance on long-tail domain-specific tasks hindered the widespread application of MLLMs in practical scenarios.
The emergence of lightweight MLLMs  [15, 16, 17, 18] has provided a good balance between parameter size and performance, alleviating the reliance on expensive computing devices and fostering the development of various downstream applications. However, there are still several challenges: (1) Most existing MLLMs use visual encoders like CLIP [13], which are trained on Internet-domain image-text data and are aligned with BERT [19, 20]. As a result, these visual encoders are not capable of covering the extensive range of visual domains and are misaligned with LLMs’ representations. (2) To adapt MLLMs to specialized domains, existing methods mainly focus on modifying the model architectures, gathering extensive related training data, or customizing the training process for the target domain. There is still no consensus framework for LLMs’ downstream adaptation. Different domains have different model designs, data formats, and training schedules.
To address these issues, there is a need for a strong visual encoder with comprehensive visual knowledge as well as a general transfer learning paradigm that allows for efficient application across downstream tasks in various domains at a low marginal cost.
In this work, we introduce Mini-InternVL, a series of powerful pocket-sized MLLMs that can be easily transferred to various specialized domains. To this end, we first enhance the representational capabilities of a lightweight visual encoder. We initialize a 300M visual encoder using the weights from CLIP and apply knowledge distillation using InternViT-6B [1] as the teacher model. Subsequently, we develop Mini-InternVL series with 1 billion, 2 billion, and 4 billion parameters, by integrating the visual encoder with the pre-trained LLMs such as Qwen2-0.5B [11], InternLM2-1.8B [9], and Phi-3-mini [21], respectively. Benefiting from the robust visual encoder, Mini-InternVL exhibit excellent multimodal performance on general multimodal benchmarks like MMBench, ChartQA, and MathVista. Remarkably, compared with InternVL2-76B, the proposed Mini-InternVL-4B achieves 90% of the performance of larger counterparts while using to 5% fewer parameters, significantly reducing computational overhead.
To further adapt our models to specific-domain downstream tasks, we introduce a straightforward yet effective transfer learning paradigm. Within this paradigm, we developed a unified transfer approach applicable to various downstream tasks, including autonomous driving, medical images, and remote sensing. This approach standardizes the model architecture, data format, and training schedule. The results demonstrate the effectiveness of this method in enhancing the model’s visual understanding and reasoning capabilities in domain-specific scenarios, enabling it to match the performance of proprietary commercial models within the target domains.
In summary, our contribution has three folds:
(1) We propose Mini-InternVL, a powerful pocket multimodal model, that not only achieves robust multimodal performance with fewer than 4 billion parameters but also easily transfers to downstream tasks across various domains at low marginal cost.
(2) We develop several design features for Mini-InternVL, including a lightweight visual encoder—InternViT-300M, that is robust for various visual domains. Additionally, we introduce a simple but effective paradigm that standardizes model architecture, data format, and training schedule for effective downstream task transfer.
(3) We comprehensively evaluate our models through extensive experimentation on general and domain-specific benchmarks. The experimental results show that our multimodal models achieve 90% of the performance using significantly fewer parameters on general multimodal benchmarks. For specific domain tasks, with minimal computational cost for fine-tuning, they can rival closed-source commercial models. We conduct a series of ablation studies to explore the impact of data sample size on domain adaptation, hoping to provide insights into the application of MLLMs in specialized domains.
2 Related Works
Multimodal Large Language Models.
Benefiting from the advancement of LLMs, the MLLMs have also achieved great progress. Early works [22, 23, 24] consider multi-modal understanding as one of the tool usage tasks and prompt the LLMs to ask other models to write a caption about the corresponding input modality so that LLMs could understand the multi-modal input. To effectively utilize the ability of pre-trained LLMs and VFMs, a series of works [25, 26, 27, 28, 1, 29, 30] propose to use a connector to align the embedding space between them, which achieve promising performance under a controllable cost. Another series of work [31, 32, 33, 7] extend pre-trained LLMs with extra layers to fuse the vision features, which reduce the number of required visual tokens inputted into LLMs while introducing extra training cost.
Recently, some works, such as Fuyu [34],MoMa [35] and Chameleon [36], propose a visual encoder-free architecture. This type of architecture consists of a single Transformer model, which is used to process both visual and textual information simultaneously without requiring an additional encoder, making it more deployment-friendly. Despite these advancements, the heavy inference cost of these MLLMs hinders their application in downstream tasks. To address such issue, a series of ligthweight MLLMs, such as MiniCPM-V [15], are proposed. However, since most of them use CLIP-L [13] as visual encoder, which is only trained on natural image domain, these models are limited to the general domain and fail to generalize to other domain. In this work, we propose InternViT-300M, which is distilled from InternViT-6B and trained on a diverse image domain.
Vision Foundation Models for MLLMs.
From a vision-centric perspective, most MLLMs utilize vision models such as CLIP [13, 37] and SigLIP [14], which are trained on large-scale web image-text data. However, such visual encoders face significant limitations in terms of parameter scale and representational ability. Several studies have explored this issue. For instance, Tong et al. [38] identified significant differences in the visual patterns of CLIP and DINOv2 [39], leading to the development of a mixture-of-features module that integrates these two VFMs. LLaVA-HR [40] introduced a dual-branch vision encoder that employs CLIP-ViT for low-resolution pathways and CLIP-ConvNext for high-resolution pathways. Similarly, DeepSeek-VL [41] utilized a dual vision encoder design, incorporating SigLIP-L for low-resolution images and SAM-B [42] for high-resolution images.
However, these methods involve excessively complex pathways, which complicates the practical application of the models. Moreover, such approaches do not resolve the issue of visual encoders lacking comprehensive visual knowledge across various domains. In contrast, InternViT [1] implements progressive image-text alignment, and acquires representational capabilities across multiple domains by performing generative training on datasets spanning various fields. We propose injecting visual knowledge from the capable vision encoder into lightweight visual models, thus avoiding the computational expense associated with iterative generative pretraining.
Domain-Specialized Adaptation of MLLMs.
Several methods have been explored to apply MLLMs to specific domains, such as GeoChat [43] and EarthGPT [44] for remote sensing, LLaVA-Med [45] and Qilin-Med-VL [46] for the medical images, ChemVLM [47] for chemistry, DriveVLM [48],DriveMLM [49] and DriveGPT4 [50] for autonomous driving. Although these methods have achieved promising results, they involve modifications to model architectures, the collection of extensive domain-specific training data, or customization of the training process for the target domain. Nonetheless, there is still no universally accepted framework for the downstream adaptation of MLLMs. we propose a straightforward yet effective transfer learning paradigm, aiming to prevent significant disparities among MLLMs in different fields that hinder interoperability.
Figure 1: Training method and architecture of Mini-InternVL. Left: We employ InternViT-6B [1] as the teacher model to perform knowledge distillation on the student model. Right: Mini-InternVL adopts the ViT-MLP-LLM architecture similar to popular MLLMs [15, 41, 27, 28, 51, 52], combining InternViT-300M with a series of lightweight LLMs through an MLP projector. Here, we employ a simple pixel shuffle to reduce the number of visual tokens to one-quarter
3 Method
In this section, we introduce Mini-InternVL, a series of lightweight multimodal large language models (MLLMs). Section 3.1 provides a comprehensive overview of Mini-InternVL. Then, Section 3.2 details InternViT-300M, a lightweight vision model developed through knowledge distillation, which inherits the strengths of a powerful vision encoder. Finally, Section 3.3 describes a transfer learning framework designed to enhance the model’s adaptation to downstream tasks.
3.1 Mini-InternVL
As shown in Figure 1, Mini-InternVL consists of three main components: InternViT, MLP Project, and LLMs. We employ InternViT-300M as our visual encoder, a lightweight vision model inherits the capabilities of a powerful vision encoder. Based on InternViT-300M, we develop three versions of Mini-InternVL: Mini-InternVL-1B, Mini-InternVL-2B, and Mini-InternVL-4B. Each version is respectively connected to the pre-trained Qwen2-0.5B [11], InternLM2-1.8B [9], and Phi-3-mini [21]. Similar to other open-source MLLMs [27, 2, 5, 52], Mini-InternVL employs an MLP projector to connect the visual encoder and the LLMs.
We adopt a dynamic resolution input strategy similar to that of InternVL 1.5 [2], which improves the model’s ability to capture fine-grained details. We also apply a pixel shuffle operation to reduce the number of visual tokens to one-quarter of the original. Consequently, in our model, a 448 448 image is represented by 256 visual tokens, enabling it to process up to 40 image crops (i.e., 4K resolution).
The training of Mini-InternVL consists of two stages: (1) Language-Image Alignment: We keep only the MLP component unfrozen during this stage. Following InternVL-1.5 [2], we use a diverse range of training datasets that encompass various tasks, including captioning, detection, grounding, and OCR. The diversity of these datasets ensures robust pre-training of Mini-InternVL, enabling the model to handle a variety of linguistic and visual elements across different tasks. (2) visual instruction tuning: We carefully select datasets to enhance the model s performance across a broad spectrum of multimodal tasks, similar to InternVL-1.5. These tasks include image captioning, chart interpretation, OCR, and cross-disciplinary reasoning. We conduct full-parameter fine-tuning with these datasets, further injecting world knowledge and teaching models to follow user instructions.
3.2 InternViT-300M
Table 1: Datasets used in knowledge distillation of visual encoder.
Type Dataset
Natural images Laion [53], COYO  [54], GRIT [29], COCO [55],
LVIS [56], Objects365 [57],
Flickr30k [58],VG [59],All-Seeing [51, 52],
MMInstruct [60],LRV-Instruction [61]
OCR TextCaps [62], Wukong-OCR [63],CTW [64],
MMC-Inst [65], LSVT [66], ST-VQA [67],
RCTW-17 [68], ReCTs [69], ArT [70],
SynthDoG [71],LaionCOCO-OCR [72],
COCO-Text [73], DocVQA [74], TextOCR [75],
LLaVAR [76],TQA [77],SynthText [78]
DocReason25K [79],Common Crawl PDF
Chart AI2D [80] ,PlotQA [81],InfoVQA [82],
ChartQA [83], MapQA [84],FigureQA [85],
IconQA [86],MMC-Instruction [87]
Multidisciplinary CLEVR-Math/Super(en) [88, 89],GeoQA+ [90],
UniChart [91],ScienceQA [92], Inter-GPS [93],
UniGeo [94],PMC-VQA [95],TabMWP [96],
MetaMathQA [97]
Other Stanford40 [98], GQA [99], MovieNet [100],
KonIQ-10k [101],ART500K [102],ViQuAE [103]
Most existing MLLMs [15, 41, 27, 28, 51, 52] employ visual encoders that are trained on web-scale image-text paired data, such as CLIP, to obtain their representations. These encoders lack comprehensive knowledge of the visual world, which needs to be acquired through iterative generative pretraining in conjunction with LLMs. Unlike other approaches that enhance the visual foundation models by using auxiliary pathways [38, 41, 16], our method directly leverages a powerful vision model that has undergone generative training on diverse datasets to transfer knowledge to a lightweight vision model. Specifically, we use InternViT-6B as the teacher model and initialize the student model s weights using CLIP-ViT-L-336px. We align the representations of the student model with those of the teacher model by computing the negative cosine similarity loss between the hidden states of the last $K$ transformer layers. The resulting model is named InternViT-300M.
The primary goal of this knowledge transfer is to inherit the pretraining knowledge embedded in InternViT-6B. To achieve this, we curate a dataset sourced from a diverse range of publicly accessible resources, as detailed in Table 1. This dataset comprises four main types of data: natural images, OCR images, charts, and multi-disciplinary images. All images are resized to a resolution of 448 $\times$448, and dynamic resolution [2] is disabled for training efficiency. Ultimately, we develop a visual encoder, termed InternViT-300M, which is infused with diverse knowledge and is adaptable to various language models.
3.3 Domain Adaptation
Figure 2: The data format of our adaptation framework. We formulate other visual tasks (Image classification, region perception, multi-view images tasks, video related tasks and visual grounding) into VQA format in our framework.
Although many studies [43, 44, 45, 50] have successfully applied MLLMs to downstream tasks, a universally accepted framework for adapting MLLMs to these applications has yet to be established. Differences in model design, data formats, and training strategies across various domains result in significant heterogeneity among MLLMs, making standardization challenging. To address this issue, we propose a straightforward yet effective transfer learning framework.
Data Format.
Instruction tuning is a crucial training stage to teach models to follow user instructions, of which the training data is formulated as visual question answering (VQA) and conversation format. VQA datasets of the downstream tasks, such as RSVQA [104] and PMC-VQA [95], are directly utilized as instruction-following data. For other conventional tasks, as shown in Figure 2, we formulate them into VQA format according to the following approaches separately:
(1) Image Classification Tasks. In most traditional classification tasks within specialized domains, a wide range of technical terms are involved. In the majority of cases, we can easily format the classification task as a multiple-choice question. Given an image <image>, the set of candidate labels $O$, and the ground truth $G\in O$, the template can be expressed as:
USER: [Image][Prompt_Prefix][Candidate Labels][Prompt_Suffix] ASSISTANT: [Ground Truth]
A direct example can be seen in our approach to remote sensing image classification, where we utilize prompts such as “Classify the image within one of the given classes: dense residential area, ..., school. Answer with one word or short phrase.”, as shown in Figure 2. This method transforms image classification tasks into multiple-choice questions. For behavior prediction of the ego vehicle in autonomous driving data, we draw inspiration from DriveLM [105] by employing templates like “Predict the behavior of the ego vehicle. Please select the correct answer from the following options: A. The ego vehicle is going straight. The ego vehicle is not moving. B. ...”.
(2) Visual Grounding Tasks. The native support for the Visual Grounding task in Mini-InternVL allows the use of a special token, <ref></ref>, to enclose the name of the object to be detected. With this token, the model can be directed to provide the object’s location enclosed within <box></box> in the format [[x1, y1, x2, y2]], where the coordinates range from 0 to 1000. This approach enables us to convert object grounding and referring expression detection into a conversational format. We extensively apply this format to remote sensing instruction data. For example, for the referring expression “1 overpass near some trees at the center”, we use “Detect <ref>1 overpass near some trees at the center</ref>” as the instruction and “<ref>1 overpass near some trees at the center</ref><box>[[x1, y1, x2, y2]]</box>” as the response.
(3) Region Perception Tasks. Region-level conversation tasks are prevalent in specialized domains. These tasks involve supplying the model with spatial location information, in addition to the question input. The model is required to focus on objects within the specified attention region to generate a response. Specifically, there are two implementation methods. The first method involves directly annotating the location on the image using bounding boxes, masks, or contours, as illustrated in Region Perception2 of Figure 2. The second method denotes the object within the question by <box>[[x1,y1,x2,y2]]</box>, where the coordinates are normalized between 0 and 1000. This notation guides the model’s attention to specific regions within the image, enabling it to perform tasks such as region-level captioning and region-specific VQA.
For example, in the context of remote sensing, the objective is to enable the model to identify the object located within the coordinates [x1, y1, x2, y2]. To achieve this, we use a prompt such as “What object is in this location<box>[[x1, y1, x2, y2]]</box>” as the input instruction, with “<ref>object name</ref><box>[[x1, y1, x2, y2]]</box>” serving as the label.
Figure 3: Qualitative Results of Mini-InternVL-DA. In the upper left corner is a multi-view image from DriveLM-nuScenes version-1.1 [105] after data processing. The color of the bounding boxes corresponds to the font color of the c tags for the objects in question (note that the input images do not contain these manually drawn bounding boxes). We show the predicted answers of our model in perception, prediction, and planning tasks. The model’s outputs align with human driving behavior.
(4) Multi-View Images. In Autonomous driving, the images are captured from six different viewpoints. As shown in Figure 3, we effectively utilize dynamic resolution to accommodate this type of data. Specifically, InternVL supports splitting images into 448$\times$448-sized tiles based on their aspect ratio. Consequently, we resize each image to 896$\times$448 pixels and then, as illustrated, combine these images in a fixed sequence, resulting in a final resolution of 2688$\times$896. This means that the images are automatically processed into 12 tiles, and an additional thumbnail is added to provide the model with global context. Furthermore, we labeled each viewpoint image with text indicating its camera position, such as “CAM_FRON”.
(5) Video Frames. InternVL supports video frames in an interleaved image format. We represent the frame sequence using a template such as “Frame1: <img><IMG_CONTEXT></img> Frame2: <img><IMG_CONTEXT></img>.”, where <IMG_CONTEXT> denotes the image tokens. For each image at a resolution of 448$\times$448, the model can accommodate sequences of up to 40 frames.
Training Strategy.
During the domain adaptation phase, we perform full-parameter fine-tuning on Mini-InternVL. For a domain-specific application scenario, we convert corresponding data into the required format and incorporate it into our training dataset. Adding a certain proportion of general multimodal data during the domain adaptation phase will not affect the performance in the specific domain, while retaining the model’s general multimodal capability. In our experiments, we find that adding general data can improve the generalization ability of the model on other tasks. Therefore, when performing domain adaptation, we can choose the appropriate general data ratio on the premise of balancing computational overhead and performance.
Table 2: Comparison with other models on multimodal benchmarks. We evaluate models using the InternVL and VLMEvalKit [106] repositories. AI2D [80], ChartQA [83], DocVQA [74], InfoVQA [82], and MMBench [107] are tested with InternVL, while MathVista [108] and OCRBench [109] use VLMEvalKit. For MMMU [110], we report scores from OpenCompass leaderboard. The Avg. Score is the average of the scores from all tested benchmarks, with the OCR-Bench score divided by 10. The values in parentheses represent the relative parameters and performance of Mini-InternVL compared to InternVL2-Llama3-76B [2], which is considered as 100%.
model open- #param MMMU MathVista AI2D ChartQA DocVQA InfoVQA OCR- MMB- MMB- Avg. Score
source (val) (testmini) Bench EN CN
GPT-4V-0409 [111] - 61.7 58.1 89.4 78.1 87.2 - 678 81.0 80.2 75.4
Gemini-Pro-1.5 [112] - 60.6 57.7 80.3 81.3 86.5 72.7 754 73.9 73.8 73.6
Claude3.5-Sonnet [113] - 65.9 67.7 94.7 90.8 95.2 - 788 79.7 80.7 81.7
GPT-4o [4] - 69.2 63.8 94.2 85.7 92.8 - 736 83.4 82.1 80.6
Cambrian-1 [114] - 50.4 53.2 79.7 75.6 75.5 - 600 81.4 - 68.0
DeepSeek-VL-1.3B [41] 2B 33.8 29.8 51.5 - - - 413 64.6 62.9 47.3
MIniCPM-V 2.0 [15] 3B 38.2 38.7 62.9 - 71.9 - 605 69.1 66.5 58.3
Qwen2-VL-2B [115] 2B 42.2 43.0 74.7 73.5 90.1 65.5 794 74.9 73.5 68.5
InternVL2-Llama3-76B [2] 76B 58.2 65.5 87.6 88.4 94.1 82.0 839 86.5 86.3 81.4
Mini-InternVL-1B 1B (1%) 36.7 37.7 64.1 72.9 81.7 50.9 754 65.4 60.7 60.6 (74%)
Mini-InternVL-2B 2B (3%) 36.3 46.3 74.1 76.2 86.9 58.9 784 73.2 70.9 66.8 (82%)
Mini-InternVL-4B 4B (5%) 48.3 58.6 78.9 81.5 89.2 67.0 788 78.6 73.9 72.8 (90%)
4 Experiments
In this section, we begin by conducting a comprehensive comparison of our Mini-InternVL with leading multi-modal large language models (MLLMs) on representative vision-language benchmarks (Section 4.1). Following this, in Section 4.2, we apply the domain adaptation framework introduced in Section 3.3 to transfer our models to three specialized domains: autonomous driving (Section 4.2.1 and Section 4.2.2), medical images (Section 4.2.3), and remote sensing (Section 4.2.4). Additionally, we perform an extensive ablation study to explore the impact of data sample size and model size on domain adaptation (Section 4.3).
4.1 Results on General Multimodal Benchmark.
Settings.
In this section, we present a comprehensive evaluation of our model’s multimodal understanding and reasoning capabilities across a variety of benchmarks. The benchmarks used in our study are categorized into four distinct types: OCR-related tasks, including DocVQA [74] OCRBench [109] and InfographicVQA [82]; chart and diagram understanding, including AI2D [80] and ChartQA [83]; general multimodal tasks, such as MMBench [107]; and multimodal reasoning, including MMMU [110] and MathVista [108]. Additionally, we report the average score of our model on the OpenCompass [116].
Results.
As shown in Table 2, Mini-InternVL demonstrates strong performance across the majority of benchmarks. Our smallest model contains only 1 billion parameters, yet it demonstrates performance comparable to 2 billion parameter models, such as DeepSeek-VL-1.3B and MiniCPM-V 2.0. Compared to other lightweight models, our Mini-InternVL-4B excels across most benchmarks, particularly in MMbench, ChartQA, DocVQA, and MathVista, where its performance is on par with commercial models like Gemini-Pro-1.5. Notably, compared to InternVL2-Llama3-76B, which utilizes the larger InternViT-6B, Mini-InternVL achieves approximately 90% of its performance while using 5% parameters. This highlights the effectiveness of our knowledge distillation strategy.
4.2 Transfer to Various Specialized Domains
4.2.1 Multi-View Image-Based Autonomous Driving
Table 3: The sources of our general datasets. In each task, we sample a specific amount of data from the data source at a predetermined ratio to balance the training dataset. In each domain adaptation case, we sample a certain amount of data from each data source.
Data Source Total Data Data Type
ShareGPT4V [117] 767k Captioning
AllSeeingV2 [52] 127k Grounding VQA
LLaVA_Instruct_zh [^1] https://huggingface.co/datasets/openbmb/llava_zh 158k VQA
DVQA [118] 200k Diagram VQA
ChartQA [119] 18k Diagram VQA
AI2D [120] 12k Diagram VQA
DocVQA [74] 10k Document VQA
GeoQA+ [121] 72k Geometric VQA
Synthdog_en [122] 30k OCR
[…]
Method #param Accuracy ChatGPT Bleu 1 Bleu 2 Bleu 3 Bleu 4 ROUGE L CIDEr Match Final Score
InternVL4Drive-v2 [123] 26B 0.7339 65.25 0.7787 0.7176 0.6608 0.6059 0.7449 0.2061 47.65 0.6002
[…]
Team NVIDIA - 0.7746 59.89 - - - - - - - 0.5884
MMFM_AD - 0.6658 63.92 - - - - - - - 0.5732
Mini-InternVL-4B 4B 0.0 54.45 0.2405 0.0801 0.0252 0.0084 0.1927 0.0018 34.30 0.3051
InternVL2-Llama3-76B 76B 0.0 52.50 0.2100 0.0884 0.0249 0.0078 0.1848 0.0001 34.22 0.2963
Mini-InternVL-DA-1B 1B 0.7007 63.84 0.7362 0.6767 0.6214 0.5678 0.7365 0.1669 39.76 0.5686
Mini-InternVL-DA-2B 2B 0.7628 65.23 0.7616 0.7012 0.6452 0.5908 0.7447 0.1914 43.24 0.5958
Mini-InternVL-DA-4B 4B 0.7296 63.97 0.7642 0.7032 0.6463 0.5914 0.7427 0.1976 42.16 0.5821
Settings.
We select DriveLM-nuScenes version 1.1 [105] as our training dataset, which contains 317k training samples and encompasses various aspects of the driving process. This dataset includes data for perception, prediction, and planning, offering a comprehensive understanding of autonomous driving scenarios.
In DriveLM-nuScenes, the images are captured from six different viewpoints. We effectively utilize dynamic resolution feature to accommodate this type of data. Specifically, Mini-InternVL supports splitting images into 448$\times$448-sized tiles based on their aspect ratio. As illustrated in Figure 3, we resize the image of each view to 896$\times$448 pixels and then combine these images in a fixed sequence, resulting in a final resolution of 2688$\times$896. This means that the images are automatically processed into 12 tiles, and an additional thumbnail is added to provide the model with global context. Furthermore, we mark the image of each view with text indicating its camera position, such as “CAM_FRON”.
As shown in Figure 3, DriveLM-nuScenes contains QA pairs with coordinates, thus we need to normalize them to a range of 0 to 1000 to align with the output of Mini-InternVL. In the dataset, objects are represented by c tags. We use a tailored prompt: “Objects are encoded using <c, CAM, [cx,cy]>, where c is the identifier, CAM indicates the camera where the object s center point is situated, and x, y represent the horizontal and vertical coordinates of the center point of the 2D bounding box.” to guide the model on the composition of c tags. The ground truth responses typically include bounding box annotations for questions requiring a list of all objects. Therefore, We annotate the bounding box as <box>[[x1,y1,y2,y3]]</box>, where <box> and </box> are special tokens in Mini-InternVL. Additionally, we incorporate general datasets to the training set, to prevent the model from losing its general domain perception capabilities, maintaining a 1:4 ratio of general to domain-specific data. The sources of the general datasets are shown in Table 3.
Finally, we conduct full-parameter fine-tuning of Mini-InternVL using 8 A100 GPUs, training the model for 1 epoch with a learning rate of 1e-5.
We report the performance of our model after transfer learning on the CVPR 2024 Autonomous Driving Challenge [105]. Furthermore, we evaluate our model on autonomous driving scenarios from MME-Realworld [124], where we separately assess its performance on Perception and Reasoning tasks.
Table 5: Results on Autonomous Driving domain of MME-RealWorld. “DA” means model after domain adaptation on DriveLM.
Method Perception Reasoning Avg.
GPT-4o [4] 21.14 26.41 24.60
Claude 3.5 Sonnet [125] 32.43 31.92 32.10
LLaVA-OneVision-7B [5] 45.77 34.08 41.75
Qwen2-VL-7B [115] 34.62 31.47 33.54
InternVL2-Llama3-76B [2] 47.46 35.71 44.30
Mini-InternVL-1B 31.34 22.47 28.96
Mini-InternVL-2B 39.86 30.13 37.25
Mini-InternVL-4B 38.96 33.11 37.39
Mini-InternVL-DA-1B 42.95 27.75 38.87
Mini-InternVL-DA-2B 50.74 40.48 47.98
Mini-InternVL-DA-4B 53.14 39.14 49.38
Results.
We test our model using DriveLM-nuScenes-version-1.1-val[105], and the results are presented in Table 4. Our final score of Mini-InternVL-2B is 0.5958, which is comparable to the best result on the CVPR 2024 Autonomous Driving Challenge Leaderboard  [^2] https://opendrivelab.com/challenge2024/#driving_with_language, InternVL4Drive-v2 [123]. Notably, our model uses only one-tenth of the parameters of InternVL4Drive-v2.
Our model scores slightly lower in the Match metric, which might be due to Mini-InternVL’s lack of proficiency in predicting object center points. InternVL4Drive-v2 [123] offers a viable solution by using Segment Anything [126] to convert object center points into object bounding boxes. In Figure 3, we show the predicted answers of our model in perception, prediction, and planning tasks, demonstrating alignment with human driving behavior. Furthermore, our 4B-parameter model performs similarly to our 2B-parameter model. We attribute this potentially to the limitations of the existing training data and evaluation criteria, which might constrain larger models from achieving significant performance gains.
As shown in Table 5, in the autonomous driving scenarios of MME-Realworld, we observe that even when using only DriveLM as domain-specific training data, our model achieves an improvement of over 10 points. The transferred model surpass the best-performing model on this task, LLaVA-OneVision-7B [5], as well as several commercial closed-source models such as GPT-4o [4] and Claude 3.5 Sonnet [125], demonstrating the strong generalization capability of our model.
4.2.2 Autonomous Driving with Temporal Information
Settings.
Using single-frame images alone is insufficient for accurate perception and prediction of vehicle behavior. Therefore, we explore temporal expansion. Specifically, we utilize instruction-following data constructed by DriveGPT4 [50] from the BDD-X dataset [127] as our training set, which comprises 26k video clips. Multiple video frames are organized as described in Section 3.3. Each training sample contains four aspects of question-answer data: Action Description, Action Justification, Speed Signal Prediction, and Turning Angle Signal Prediction. We set the proportion of general to domain-specific data at 1:1.
Following DriveGPT4 [50], we report several metric scores widely used in the NLP community, including CIDEr, BLEU4, and ROUGE-L, to evaluate the action descriptions and justifications. For open-loop control signal prediction, we use root mean squared error (RMSE) and threshold accuracies ($A_{\tau}$) for evaluation.
Table 6: The results on action tasks of BDD-X dataset. We provide evaluation results on action description, action justification, and full-text generation(i.e.,combining description and justification). B4 stands for BLEU4. “DA” means model after domain adaptation on BDD-X.
Method Description Justification Full
CIDEr B4 ROUGE CIDEr B4 ROUGE CIDEr B4 ROUGE
ADAPT [128] 219.35 33.42 61.83 94.62 9.95 32.01 93.66 17.76 44.32
DriveGPT4 [50] 254.62 35.99 63.97 101.55 10.84 31.91 102.71 19.00 45.10
Mini-InternVL-DA-1B 223.85 34.17 62.11 95.52 9.70 32.58 83.72 16.78 44.29
Mini-InternVL-DA-2B 242.14 35.77 63.03 105.06 10.63 32.46 98.47 18.05 44.52
Mini-InternVL-DA-4B 237.41 35.94 63.67 104.62 9.51 32.23 97.42 17.70 44.98
Table 7: Quantitative results of control signals prediction on BDD-X test dataset. RMSE denotes the root mean squared error, and $A_{\tau}$ measures the proportion of test samples with prediction errors less than $\tau$. “DA” means model after domain adaptation on BDD-X.
Method Speed(m/s) Turningangle(degree)
RMSE $A_{0.1}$ $A_{0.5}$ $A_{1.0}$ $A_{5.0}$ RMSE $A_{0.1}$ $A_{0.5}$ $A_{1.0}$ $A_{5.0}$
ADAPT [128] 3.02 9.56 24.77 37.07 90.39 11.98 27.93 66.83 75.13 89.45
DriveGPT4 [50] 1.30 30.09 60.88 79.92 98.44 8.98 59.23 72.89 79.59 95.32
Mini-InternVL-DA-1B 1.28 29.44 60.38 79.34 98.67 9.45 59.34 73.54 80.28 92.76
Mini-InternVL-DA-2B 1.26 27.96 59.23 80.06 98.78 9.52 57.40 72.54 80.06 92.04
Mini-InternVL-DA-4B 1.31 28.84 60.94 78.78 98.61 9.46 59.12 73.15 80.17 92.65
Results.
We report our scores on the BDD-X testing set in Table 6 and Table 7. Although our model has not undergone pre-training on large amounts of proprietary domain data like DriveGPT, it still performs comparably to DriveGPT4 across the four tasks and surpasses ADAPT [128]. Note that the performance of the three models on this dataset is similar, which aligns with the observations discussed in Section 4.2.1.
4.2.3 Medical Image Question Answering
Table 8: The sources of our medical data. The table presents the data types and the sample sizes collected from each source. We sampled a total of 500k image-text pairs from multiple publicly available datasets of different data types as our medical training data.
Data Size Description
PMC-VOA [129] 238k
MedICaT [130] 31k The datasets include image-text pairs containing
PMC-Image [129, 95, 131] 29k X-rays, pathology images, and images of affected areas,
Open-i [132] 1k extracted from open-source websites or journal articles.
MedPix [133] 6k
Quilt-1M [134] 95k A medical dataset includes image-text pairs of
histopathology images.
RP3D [131] 82k A medical dataset includes image-text pairs of X-ray images.
MIMIC-CXR [135] 14k
Retina Image Bank [136] 4k A medical dataset includes image-text pairs of retinal images.
Table 9: The results of our model on GMAI-MMBench. The results of other models are taken from GMAI-MMBench Leaderboard. “DA” means model after domain adaptation on medical data. After supervised fine-tuning, our model shows significant improvement and outperforms several medical-specialized models (e.g., LLaVA-Med, RadFM) and some commercial closed-source models (e.g., Claude3-Opus) on most metrics.
Model Size Seg C Seg M 2D Cls 2D Det 2D Mcls_acc 2D Mcls_recall
Qwen-VL-Chat [28] 9.6B 34.45 35.20 39.55 22.04 42.88 81.23
LLaVA-NeXT-mistral-7B [137] 7.6B 36.29 35.20 39.34 27.87 44.05 47.70
LLaVA-Med [45] - 18.45 18.97 21.15 17.14 45.84 41.19
RadFM [131] 14B 20.43 20.27 25.71 18.83 40.98 57.45
Claude3-Opus [125] - 33.56 33.36 32.17 24.72 45.31 38.98
GPT-4V [3] - 47.87 46.58 42.24 30.32 45.21 40.59
Mini-InternVL-1B 1B 34.30 34.55 36.02 24.08 21.67 8.57
Mini-InternVL-2B 2B 35.33 35.61 38.08 25.31 43.52 16.13
Mini-InternVL-4B 4B 36.60 36.99 38.74 26.01 43.99 16.25
Mini-InternVL-DA-1B 1B 38.67 39.44 35.87 23.09 22.79 8.99
Mini-InternVL-DA-2B 2B 40.22 39.46 39.34 25.59 44.33 16.20
Mini-InternVL-DA-4B 4B 41.41 40.45 41.34 24.84 44.33 16.59
Settings.
We utilize several publicly available medical image-text datasets to improve the model’s understanding of medical images. These datasets include a wide range of medical images, such as photos, X-rays, and pathology images. The datasets include PMC-OA [129], MedICaT [130], PMC-Image [129, 95, 131], Open-i [132], MedPix [133], Quilt-1M [134], RP3D [131], MIMIC-CXR [135], and Retina Image Bank [136], which together provide a substantial collection of medical images. From these datasets, we sampled 500k image-text pairs for the model’s training set. Finally, We add general data in a 1:1 ratio to the domain-specific data and conduct full-parameter training of the model for one epoch.
Results.
In this section, we present the performance of Mini-InternVL and its fine-tuned variant, Mini-InternVL-DA, on a comprehensive medical AI benchmark, GMAI-MMBench [138]. Our evaluation is conducted using the VLMEvalKit [^3] https://github.com/open-compass/VLMEvalKit framework. Table 9 illustrates the performance of various models on the medical VQA tasks.
After supervised fine-tuning, our model shows significant improvement across most evaluation metrics. Specifically, it excels in 2D classification (2D Cls), 2D detection (2D Det), and 2D multi-class accuracy (2D Mcls_acc). These results highlight its strong multimodal understanding capabilities in complex medical visual question-answering tasks.
Furthermore, our model of 4B size outperforms several medical-specialized models (e.g., LLaVA-Med [45], RadFM [131]) and some commercial closed-source models (e.g., Claude3-Opus [125]) on most metrics. However, there is no improvement in multiple-choice questions after SFT, which we attribute to the lack of multiple-choice question data in the training dataset.
4.2.4 Remote Sensing
Settings.
The training data is summarized in Table 10. The GeoChat instruction set [43] serves as the primary component of our training dataset. To enrich the dataset with high-resolution imagery, we also incorporate the RSVQA-HR dataset [104]. Additionally, we include 100k VQA instances sampled from FIT-RS [139] to further expand our training set. For the visual grounding task, we integrate the DIOR-RSVG dataset [140] into our training process. All data are reformatted according to the methods outlined in Section 3.3.
A single epoch of training on the visual grounding data is found to be insufficient, so we repeat the DIOR-RSVG training samples multiple times. Finally, we incorporate 20% of the total general domain training samples into the training data and conduct training following the settings described in Section 4.2.1.
We assess the performance of our transferred model using the RSVQA dataset for the VQA task and the DIOR-RSVG dataset for the visual grounding task. Following the methodology outlined in [43], we chose the Presence, Comparison, and Rural/Urban subsets of the RSVQA-LR and RSVQA-HR datasets for assessment.
Table 10: Details on the training samples used to adapt remote sensing.
Dataset Data type Size
GeoChat [43] Detailed Description 75k
Multi-Round Conversation 65k
Complex Questions 10k
VQA 91.5k
Region Captioning 40k
Visual Grounding 25k
RSVQA-HR [104] VQA 50k
FIT-RS-VQA [139] VQA 100k
DIOR-RSVG [140] Visual Grounding 26k
Table 11: The results on remote sensing VQA and visual Grounding tasks. For the VQA datasets,we omit area and count questions during evaluation. “DA” means model after domain adaptation on remote sensing.
Method Size RSVQA-LR RSVQA-HR-Test1 RSVQA-HR-Test2  DIOR-RSVG
Rural/Urban Presence Compare Avg. Presence Compare Avg. Presence Compare Avg. (acc@0.5)
RSVQA [141] - 90.00 87.46 81.50 86.32 90.43 88.19 83.12 86.26 85.94 77.50 -
Bi-Modal [142] - 92.66 91.06 92.66 91.63 92.03 91.83 84.98 89.37 89.62 80.54 -
EasyToHard [143] - 91.67 90.66 87.49 89.94 91.39 89.75 93.97 87.97 87.68 79.06 -
GeoChat [43] 7B 94.00 91.09 90.33 90.70 - - - - 58.45 83.19 72.30
SkyEyeGPT [144] - 75.00 88.93 88.63 84.19 84.95 85.63 85.29 83.50 80.28 81.89 88.59
SkySenseGPT [139] - 95.00 91.07 92.00 92.69 - - - 69.14 84.14 76.64 -
Mini-InternVL-4B 4B 66.00 64.64 73.26 69.55 62.42 79.20 71.72 65.73 79.70 73.55 16.89
InternVL2-Llama3-76B 76B 61.00 66.29 79.61 73.77 60.79 77.47 70.04 63.30 78.32 71.71 29.65
Mini-InternVL-DA-1B 1B 95.00 81.39 91.6 87.37 92.24 91.76 91.98 89.38 90.91 90.24 89.73
Mini-InternVL-DA-2B 2B 93.00 87.07 91.85 89.87 92.33 92.21 92.27 89.60 90.86 90.30 89.24
Mini-InternVL-DA-4B 4B 92.00 85.69 92.18 89.46 92.42 92.12 92.25 89.25 90.92 90.18 92.04
Results.
Table 11 presents the performance of our models on remote sensing VQA and visual grounding tasks. On the RSVQA task, our model demonstrates strong performance under both high-resolution and low-resolution conditions. Unlike existing remote sensing MLLMs such as GeoChat [43] and SkySenseGPT [144], which support only single-resolution images, our model leverages dynamic resolution to effectively benefit from high-resolution training data. Compared to traditional models in the remote sensing domain such as RSVQA [141], Bi-Modal [142], and EasyToHard [143] our model achieves superior scores on both RSVQA-HR-Test1 and RSVQA-HR-Test2, showcasing its generalization ability. Furthermore, our models of three different sizes outperform SkyEyeGPT on DIOR-RSVG, indicating that our framework can effectively model visual grounding tasks.
4.3 Ablation Study
Table 12: Comparison between Mini-InternVL-2B with Mini-InternVL-CLIP-2B across various tasks. We test the model on DriveLM val set after finetuning them using domain-specific data.
Method MMB- MMB- ChartQA DocVQA InfoVQA MMMU MME-RW DriveLM
EN CN (test) (val) (val) (AD)
Mini-InternVL-CLIP-2B 70.3 68.1 70.9 77.5 49.6 32.9 43.7 0.580
Mini-InternVL-2B 73.2 70.9 76.2 85.9 57.7 34.3 48.0 0.578
(a) The impact of the proportion of general data.
(b) The impact of the number of training samples.
Figure 4: The impact of the proportion of general data and the number of training samples. MME-RW-AD, MMB, and DriveLM refer to the Autonomous Driving domain of MME-RealWorld, the general multimodal benchmark MMBench, and DriveLM Challenge, respectively.
Figure 5: Performance of 3 methods at varying training steps on 4 benchmarks. After 1500 steps, the model’s performance converges. The scores for the three adaptation methods remain stable across different training steps on the general multimodal benchmark MMBench [107].
Table 13: Comparing the per-GPU memory consumption and training speed of the 3 methods. Our experiment use ZeRO1 [145] and is conducted on 8 A100 GPUs.
Method GPU memory Training Speed
(GB) (iterations per hour)
LoRA 23.40 263.02
Freezing ViT 29.87 260.76
Full-parameter 33.11 185.13
The Importance of Knowledge Distillation.
We demonstrate the effectiveness of knowledge distillation in this study. We train a new 2B-sized MLLM, which we refer to as Mini-InternVL-CLIP-2B. Specifically, we maintain the structure of the MLLM while replacing the visual encoder from InternVIT with CLIP-Large-336 [13], employing the same training recipe as described in Section 3.1. As shown in Table 12, the results across multiple benchmarks indicate that Mini-InternVL-CLIP-2B significantly outperforms Mini-InternVL-CLIP-2B on general benchmarks, particularly in document-related tasks. This clearly illustrates that our knowledge distillation in our method effectively enables InternVIT to acquire visual world knowledge. Furthermore, we adapt Mini-InternVL-CLIP-2B to autonomous driving tasks, and the results demonstrate the advantages of our model in proprietary domain transfer.
The Impact of Data Ratio Balancing.
In this study, we investigate the effect of the proportion between general data and domain-specific data on model transferability. We conducted experiments in the autonomous driving domain, utilizing all samples from DriveLM and supplementing them with general training data at multiples of $r$ times the number of DriveLM samples. The results are shown in Figure 4(a). Our findings indicate that relying exclusively on domain-specific training data does not yield optimal performance on downstream tasks. Introducing a specific ratio of general data not only enhances performance on domain-specific tasks but also reduces performance degradation on general multimodal benchmarks. This demonstrates that incorporating an appropriate proportion of general data is crucial for improving the model’s generalization ability and maintaining its general capabilities. For our autonomous driving scenario, we observe that performance peaks at $r=0.25$; beyond this point, performance slightly declines as $r$ increases. This suggests that we can achieve benefits without significantly increasing computational load.
Influence of Training Sample Size.
We investigate how varying the quantity of training samples affects performance on downstream tasks. In this experiment, we use different amounts of training data while maintaining a 1:1 ratio between general data and domain-specific data, as shown in Figure 4(b) Training the model with only one-quarter of the full dataset significantly reduces computational load during training while resulting in only a minor loss in performance. Notably, the model’s score on general benchmarks remains largely unchanged when the proportion of different training data is kept constant.
Effect of Different Adaptation Methods.
In this study, we examine the effects of different adaptation methods LoRA, freezing the visual encoder, and full-parameter fine-tuning on model performance. Using the dataset described in  Section 4.2.4, we apply these three methods to train the model across varying numbers of steps and evaluate its performance on four benchmarks: general multimodal VQA, remote sensing VQA, and visual grounding. We record memory consumption and training speed during the process, as shown in  Table 13. As illustrated in  Figure 5, model performance converges after 1500 steps, with each method exhibiting distinct performance ceilings. Notably, full-parameter fine-tuning achieves the highest scores on domain-specific tasks. Additionally, we find that LoRA underperforms on the visual grounding task, and freezing the vision encoder strikes a balance between performance and computational efficiency. The scores for all three adaptation methods remain stable across different training steps on the general multimodal benchmark, maintaining strong performance in general domain even with extended training.
5 Conclusion
In this work, we introduce Mini-InternVL, a series of lightweight, open-source MLLMs designed to tackle the challenges of deploying MLLMs in resource-constrained environments. Mini-InternVL utilizes InternViT-300M as a compact visual encoder, integrating world knowledge across multiple domains through knowledge distillation from a more capable teacher model, thereby addressing the limitations of encoders like CLIP-ViT. Mini-InternVL achieves approximately 90% of the performance of larger models using significantly fewer parameters,excelling particularly in tasks such as OCR and domain-specific image understanding. To facilitate the application of small-scale multimodal models in specialized fields, we employe a unified transfer format, enabling our models to be effectively adapted to multiple specific domains, where they achieve comparable performance to other domain-specific approaches. We hope that this work provides valuable insights into the application of MLLM.